# Stochastic Data-Driven Predictive Control: Chance-Constraint Satisfaction with Identified Multi-step Predictors

### v1

@@ VI Numerical Example
In the following, we illustrate the proposed framework using a mass-spring-damper chain. Computations were conducted in Python on a server instance with 16 GB RAM and 8-core allocation from an AMD EPYC 9654 96-core processor. The optimization problems are solved via ECOS [28] through the CVXPY interface [29]. The numerical example’s code is accessible online at: https://github.com/haldunbalim/MultistepPredictors.
@@
In Figure 1, we show two-sided probabilistic reachable sets for the last mass’s position measurements with probability level $p=0.9$. The sequential propagation results in overly conservative estimates while the multi-step predictors closely approximate the ground truth. Furthermore, the proposed strategy is less conservative than the ellipsoidal set-based approach, since we directly use distribution of parameters. Finally, the small difference between our approach and the sampling-based estimate of the probabilistic reachable set highlights the minimal conservatism of the proposed approach.
@@
In the following, we highlight the ability of the proposed framework to enforce the chance constraints (2b). We simulate the system and estimate state-space and multi-step models for $10^{3}$ times. The initial state is set with a mean $\bar{x}_{0}\equiv-0.2$ and a covariance matching the steady-state derived from the true system parameters. Input constraints are set as $|u_{t}|\leq 2.5$ and chance constraints on position measurements are set as $\Pr[y_{t,i}\leq 0.05]\geq p=0.9$. We consider the cost matrices $Q_{\mathrm{c}}=I$ and $R_{\mathrm{c}}=10^{-1}I$. Table I shows the number of instances where the considered methods find feasible solutions. As shown, the multi-step methods significantly outperform the sequential method in this example, in accordance with Fig. 1. Moreover, the proposed constraint tightening strategy yields feasible solutions in significantly more cases compared to the strategy using confidence ellipsoids.
TABLE I: Number of instances where the predictive control problem is feasible and the total offline computation time (including parameter identification for state-space model) required for each method.
@@
Finally, we compare the computational time, and constraint satisfaction of the proposed scheme against the stochastic MPC (6) with true system parameters. Out of $10^{3}$ trials, we neglect $8$ instances where problem (16) was infeasible. We compute $f_{\epsilon,k,j}$ in Eq. 17 using Davies’ method [30] and line-search and approximate it using sampling if numerical issues arise. Results for this experiment are presented in Table II. As shown, the proposed strategy successfully enforces the desired chance constraints.
@@
Exact model knowledge (6) 5.1 10.0
Proposed (data-driven) (16) 40.4 3.7
@@ VII Conclusion
We presented a data-driven framework for addressing the stochastic optimal control problem using multi-step predictors. We first propose an identification and uncertainty quantification scheme for multi-step predictors utilizing an intermediate state-space identification step. The identified multi-step predictors are used to directly propagate uncertainty in the parameter estimates, allowing for minimally conservative constraint tightening strategies that ensure chance constraint satisfaction. Finally, we provided a numerical example, highlighting the ability to guarantee satisfaction of probabilistic constraints while significantly reducing conservatism compared to state-of-the-art approaches. Future work will focus on developing a framework that addresses the receding horizon case with closed-loop guarantees, by incorporating ideas from using ideas from [9, 17].

### v2

@@ VI Numerical Example
In the following, we illustrate the proposed framework using a mass-spring-damper chain. Computations were conducted in Python on a server instance with 16 GB RAM and 8-core allocation from an AMD EPYC 9654 96-core processor. The optimization problems are solved via ECOS [25]. The numerical example’s code is accessible online at: https://github.com/haldunbalim/MultistepPredictors.
@@
In Figure 1, we show two-sided probabilistic reachable sets for the last mass’s position measurements with probability level $p=0.9$. The sequential propagation results in overly conservative estimates while the multi-step predictors closely approximate the ground truth. Furthermore, the proposed strategy is less conservative than the ellipsoidal set-based approach, since we directly use distribution of parameters. Finally, the small difference between our approach and the sampling-based estimate of the probabilistic reachable set highlights the reduced conservatism of our approach.
@@
In the following, we highlight the ability of the proposed framework to enforce the chance constraints (2b). We simulate the system and estimate state-space and multi-step models for $10^{3}$ times. The initial state is set with a mean $\bar{x}_{0}\equiv-0.2$ and a covariance matching the steady-state derived from the true system parameters. Input constraints are set as $|u_{t}|\leq 2.5$ and chance constraints on position measurements are set as $\Pr[y_{t,i}\leq 0.05]\geq p=0.9$. We consider the cost matrices $Q_{\mathrm{c}}=I$ and $R_{\mathrm{c}}=10^{-1}I$. Table I shows the number of instances where the considered methods find feasible solutions and offline design time including state-space identification. As shown, the multi-step methods significantly outperform the sequential method in this example, in accordance with Fig. 1. Moreover, the proposed constraint tightening strategy yields feasible solutions in significantly more cases compared to the strategy using confidence ellipsoids.
TABLE I: Number of instances where the control problem is feasible and offline computation time required for each method.
@@
Finally, we compare the computational time, and constraint satisfaction of the proposed scheme against the stochastic MPC (6) with true system parameters. Out of $10^{3}$ trials, we neglect $8$ instances where problem (16) was infeasible. We compute $f_{\epsilon,k,j}$ in Eq. 17 using Davies’ method [24] and approximate it using sampling if numerical issues arise. The increased computational time of the proposed approach (16) is due to the second order cone constraints that address the parametric uncertainty. Results for this experiment are presented in Table II. As shown, the proposed strategy successfully enforces the desired chance constraints. The proposed approach achieves constraint violations below $10\%$, highlighting the guarantees of Theorem 1 and its conservatism relative to the idealized solution.
@@
Exact model (idealized) (6) 5.1 10.0
Proposed (robust) (16) 40.4 3.7
@@ VII Conclusion
We presented a data-driven framework for addressing the stochastic optimal control problem using multi-step predictors. We first propose an identification and uncertainty quantification scheme for multi-step predictors utilizing an intermediate state-space identification step. The identified multi-step predictors are used to directly propagate uncertainty in the parameter estimates, allowing for simple constraint tightening strategies that ensure chance constraint satisfaction. Finally, we provided a numerical example, highlighting the ability to guarantee satisfaction of probabilistic constraints while significantly reducing conservatism compared to state-of-the-art approaches. Future work will focus on developing a framework that addresses the receding horizon case with closed-loop guarantees, by incorporating ideas from [9, 16]. Extending the proposed techniques to nonlinear and unstable systems is an interesting avenue for research.